OpenStudy (anonymous):

x = 0.5y + w2 Solve for y.

OpenStudy (anonymous):

@Jack1

OpenStudy (anonymous):

I came up with 2x – w2 = y

OpenStudy (yellowlegoguy99):

Boring

ganeshie8 (ganeshie8):

careful, first step is subtracting w2 here

OpenStudy (anonymous):

Oh

ganeshie8 (ganeshie8):

x = 0.5y + w2 subtract w2 both sides x - w2 = 0.5y

OpenStudy (phi):

the next step is multiply both sides (and *all* terms) by 2

OpenStudy (anonymous):

I came up with 2(x – w2) = y

OpenStudy (anonymous):

Yes I know @phi

OpenStudy (phi):

yes, looks good. you could distribute the 2, if you are doing multiple choice and your answer does not match any of the choices

OpenStudy (anonymous):

Is 2(x – w2) = y the correct answer?

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark} \)

OpenStudy (anonymous):

It's not multiple choice it's a word problem, I had to create my own equation and solve for a different variable.

ganeshie8 (ganeshie8):

^ distribute 2 if u want

OpenStudy (anonymous):

Thanks a lot both of you.

ganeshie8 (ganeshie8):

np :)